OpenStudy (anonymous):

Find three consecutive odd integers that when the 1st is added to twice the second and three time the 3rd, the result is 58

OpenStudy (anonymous):

I'm not from us, odd is divisible by two or not divisible by 2?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Let x be the first odd number then the other two consecutive odd integers will be (x+2) and (x+4)

OpenStudy (anonymous):

Now, x+2(x+2) + 3(x+4)=58

OpenStudy (anonymous):

SOLVE FOR X

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

yes thank you.

OpenStudy (anonymous):

WELCOME

OpenStudy (anonymous):

is the first x+2 supppsed to be 2x?

OpenStudy (anonymous):

Nope...... its x+2

OpenStudy (anonymous):

how do I solve then?

OpenStudy (anonymous):

x+2(x+2) + 3(x+4)=58 x+2x+4+3x+12=58 6x+16=58 6x=42 x=7

OpenStudy (anonymous):

7, 9, 11

OpenStudy (anonymous):

YEP

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome